OpenStudy (anonymous):

Write the equation of the line that passes through (–4, 3) and is perpendicular to the line y= 7/5x-3/5

OpenStudy (anonymous):

perpendicular lines have "flipped" values and their signs are changed. For example, if slope 1 = (2/3), a perpendicular slop would have to be (-3/2)

OpenStudy (anonymous):

The slope in y= (7/5)x - (3/5) is (7/5). Which means it's perpendicular slope is...

OpenStudy (anonymous):

-5/7

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so how do i write it in standard form?

OpenStudy (anonymous):

So now you use point-intercept form, using the point (-4,3) with slope (-5/7)

OpenStudy (anonymous):

y-y1=m(x-x1), where -4 is x1, 3 is y1, and m=(-5/7)

OpenStudy (anonymous):

Whats the equation for point intercept form?

OpenStudy (anonymous):

y-3=(-5/7)(x+4)

OpenStudy (anonymous):

y-(3)=(-5/7)(x-(-4)) y-3=(-5/7)x-(20/7) You distribute the (-5/7) to the x and the 4. Remember minus a negative makes it positive. y-3=(-5/7)x-(20/7) solve for y

OpenStudy (anonymous):

y-3=(-5/7)x-(20/7) +3 +3 y=(-5/7)x+(1/7)

OpenStudy (anonymous):

Now I'm going to check to make sure it is right

OpenStudy (anonymous):

It's correct.